# inkscapelatex=false

### v1

@@ 5.2.1   Software-based and Compiler-based preemptive scheduling
Relies on system software to manage task interruptions, providing flexibility for preemption without requiring specific hardware support. This approach is seen in frameworks like EffiSha introduced by Chen et al. [58], which enable efficient preemptive scheduling for GPU tasks. While it allows for easier integration with existing hardware, the reliance on software can introduce overhead due to frequent context switching, which may affect performance. Table LABEL:tab:soft provides an overview of preemptive scheduling methods (software-based and compiler-based). A detailed discussion of these methods is provided in the book.
@@ 5.2.2   OS-based preemptive scheduling
In OS-based preemptive scheduling, preemption is integrated directly into the operating system, which manages hardware resources more efficiently. Operating systems like VxWorks introduced by Barbalace et al. [59] and RTLinux by Ip et al. [60] have built-in support for preemptive scheduling, allowing them to handle both hard and soft real-time tasks with minimal delay. By coordinating hardware and software layers, OS-based approaches minimize latency, making them suitable for systems requiring real-time guarantees. Table LABEL:tab:os provides an overview of preemptive scheduling methods (OS-based). A detailed discussion of these methods is provided in the book.
@@ 5.2.3   Hardware-based preemptive scheduling
This scheduling relies on specific hardware mechanisms to support preemption. This reduces the overhead of task switching and enables faster responses to real-time events. Systems like those proposed by Lin et al. [61] use hardware features for lightweight context switching, ensuring that high-priority tasks meet strict deadlines while maintaining system performance. Hardware-based solutions are especially useful in latency-sensitive environments where responsiveness is critical. Table LABEL:hw provides an overview of preemptive scheduling methods (hardware-based). A detailed discussion of these methods is provided in the book.
@@ 6   Implicit Real-Time Scheduling
In the context of GPU scheduling, implicit real-time scheduling refers to the use of conventional scheduling methods that aren’t specifically designed for real-time applications but are capable of supporting real-time tasks under certain conditions. These methods typically rely on kernel concurrency and the inherent characteristics of the scheduler, such as fixed-priority or round-robin mechanisms, to meet deadlines indirectly. Although they are not tailored for the strict requirements of real-time systems, in scenarios where timing demands are moderately flexible, these schedulers can achieve satisfactory performance by using their natural scheduling behavior. This approach is generally less complex and can be effective in systems where real-time support is beneficial but not critical. Table LABEL:tab:merged_table presents an overview of implicit real-time methods; for a detailed analysis of these works, readers are encouraged to consult the referenced book.
@@ 7   Future Directions and Applications
The interplay of GPUs with real-time systems has driven notable advancements across diverse domains that require both high computational throughput and stringent latency guarantees. In robotics, real-time systems are foundational to essential functions such as localization, collision avoidance, and sensor data processing [98, 99]. Similarly, embedded systems exploit real-time capabilities to ensure dependable operation in automotive control and industrial automation applications [100, 101]. The adoption of GPU-accelerated real-time processing in medical imaging has significantly improved the efficiency and responsiveness of data-intensive tasks [102]. Additionally, telecommunications systems benefit from real-time frameworks that enhance data transmission and Quality-of-Service management [103]. In cloud computing and the Internet of Things (IoT), real-time processing enables effective monitoring and control, supporting scalable and responsive infrastructures. Autonomous systems, including vehicles and drones, rely extensively on real-time decision-making mechanisms for navigation and obstacle avoidance. These topics will be explored in greater detail throughout this book.
@@ 8   Conclusion
GPUs have rapidly evolved into indispensable parallel computing platforms that significantly accelerate complex workloads in real-time and embedded systems, especially in critical domains such as autonomous vehicles and machine learning. Despite their computational power, inherent architectural features, such as non-preemptive execution, variable kernel latencies, and resource contention, present substantial challenges in meeting the strict timing and predictability demands of real-time applications. While recent research has made notable progress through advanced scheduling algorithms, multitasking frameworks, and resource management techniques to improve GPU responsiveness and utilization, important challenges remain. These include implementing efficient preemption mechanisms and reducing overheads without requiring extensive code modifications. Therefore, future efforts must focus on developing comprehensive, adaptive strategies that effectively balance performance, predictability, and energy efficiency to enable the reliable integration of GPUs in safety-critical and latency-sensitive real-time systems.

### v2

@@ 4.2.1  Software-based and Compiler-based preemptive scheduling
Relies on system software to manage task interruptions, providing flexibility for preemption without requiring specific hardware support. This approach is seen in frameworks like EffiSha introduced by Chen et al. [chen2017effisha], which enable efficient preemptive scheduling for GPU tasks. While it allows for easier integration with existing hardware, the reliance on software can introduce overhead due to frequent context switching, which may affect performance. Table LABEL:tab:soft provides an overview of preemptive scheduling methods (software-based and compiler-based). A detailed discussion of these methods is provided in the book.
@@ 4.2.2  OS-based preemptive scheduling
In OS-based preemptive scheduling, preemption is integrated directly into the operating system, which manages hardware resources more efficiently. Operating systems like VxWorks introduced by Barbalace et al. [barbalace2008performance] and RTLinux by Ip et al. [ip2001performance] have built-in support for preemptive scheduling, allowing them to handle both hard and soft real-time tasks with minimal delay. By coordinating hardware and software layers, OS-based approaches minimize latency, making them suitable for systems requiring real-time guarantees. Table LABEL:tab:os provides an overview of preemptive scheduling methods (OS-based). A detailed discussion of these methods is provided in the book.
@@ 4.2.3  Hardware-based preemptive scheduling
This scheduling relies on specific hardware mechanisms to support preemption. This reduces the overhead of task switching and enables faster responses to real-time events. Systems like those proposed by Lin et al. [Lin2016] use hardware features for lightweight context switching, ensuring that high-priority tasks meet strict deadlines while maintaining system performance. Hardware-based solutions are especially useful in latency-sensitive environments where responsiveness is critical. Table LABEL:hw provides an overview of preemptive scheduling methods (hardware-based). A detailed discussion of these methods is provided in the book.
@@ 5  Implicit Real-Time Scheduling
Implicit real-time scheduling refers to the use of conventional scheduling methods, which aren’t specifically designed for real-time applications but can still support real-time tasks under certain conditions. These methods typically rely on kernel concurrency and the inherent characteristics of the scheduler, such as fixed-priority or round-robin scheduling, or other simpler scheduling techniques, to meet deadlines indirectly. Although they are not tailored for the strict requirements of real-time systems, in scenarios where timing demands are moderately flexible, these schedulers can achieve satisfactory performance by using their natural scheduling behavior. This approach is generally less complex and can be effective in systems where real-time support is beneficial but not critical. Table LABEL:tab:merged_table presents an overview of implicit real-time methods; for a detailed analysis of these works, readers are encouraged to consult the referenced book.
@@ 6  Future Directions and Applications
The interplay of GPUs with real-time systems has driven notable advancements across diverse domains that require both high computational throughput and stringent latency guarantees. In robotics, real-time systems are foundational to essential functions such as localization, collision avoidance, and sensor data processing [matsuoka2011robotics, laurent2014embedded]. Similarly, embedded systems exploit real-time capabilities to ensure dependable operation in automotive control and industrial automation applications [lee2007handbook, verma2015slip]. The adoption of GPU-accelerated real-time processing in medical imaging has significantly improved the efficiency and responsiveness of data-intensive tasks [eklund2013medical]. Additionally, telecommunications systems benefit from real-time frameworks that enhance data transmission and Quality-of-Service management [atdelzater2000qos]. In cloud computing and the Internet of Things (IoT), real-time processing enables effective monitoring and control, supporting scalable and responsive infrastructures. Autonomous systems, including vehicles and drones, rely extensively on real-time decision-making mechanisms for navigation and obstacle avoidance. These topics will be explored in greater detail throughout this book.
@@ 7  Conclusion
GPUs have rapidly evolved into indispensable parallel computing platforms that significantly accelerate complex workloads in real-time and embedded systems, especially in critical domains such as autonomous vehicles and machine learning. Despite their computational power, inherent architectural features, such as non-preemptive execution, variable kernel latencies, and resource contention, present substantial challenges in meeting the strict timing and predictability demands of real-time applications. While recent research has made notable progress through advanced scheduling algorithms, multitasking frameworks, and resource management techniques to improve GPU responsiveness and utilization, important challenges remain. Future work must address the challenge of implementing efficient preemption mechanisms and reducing their associated overheads, while also adapting to emerging hardware features that increasingly support finer-grained preemption. Consequently, future efforts should focus on developing comprehensive and adaptive strategies that balance performance, predictability, and energy efficiency to enable the reliable integration of GPUs in safety-critical and latency-sensitive real-time systems.